OpenStudy (anonymous):

Two angles are complementary. The sum of the measure of the first angle and one-fourth the second angle is 57.75 degrees. Find the measures of the angles.

hero (hero):

I've seen this before

hero (hero):

x + y = 90 x + y/4 = 57.75

hero (hero):

Systems of equations

jimthompson5910 (jim_thompson5910):

Let x and y be the two angles. Since they are complementary, they add to 90 degrees. So x + y = 90. Solve for y to get y = 90-x Because "The sum of the measure of the first angle and 1/4 of the second angle is 57.75", we know that x+y/4 = 57.75 Multiply every term by 4 to get 4x+y=231 4x+y=231 4x+90-x=231 4x-x=231-90 3x=141 x=141/3 x=47 y = 90-x y = 90-47 y = 43 So the smaller angle is 43 degrees while the other angle is 47 degrees.